Find the complementary and supplementary angles of 45deg Express your answers in radians.

and

Rewrite -11pi over 4 in degrees
thank you

Answer by cleomenius(959)

The complement would be 90 degrees - 45 degrees = 45 degrees.
The supplement would be 180 degrees - 45 degrees = 135 degrees.
To convert to radians
45 * pi/180 = pi/4 radians
135 * pi/180 = 3pi/4 radians
-11pi/4 * 180/pi = -495 degrees
Cleomenius.